# Spline Dimensional Decomposition with Interpolation-based Optimal Knot Selection for Stochastic Dynamic Analysis

### v1

@@ Abstract
Forward uncertainty quantification in dynamic systems is challenging due to non-smooth or locally oscillating nonlinear behaviors. Spline dimensional decomposition (SDD) effectively addresses such nonlinearity by partitioning input coordinates via knot placement, yet its accuracy is highly sensitive to the location of internal knots. We propose a computationally efficient, interpolation-based method for optimal knot selection in SDD. The method involves three steps: (1) interpolating input-output profiles, (2) defining subinterval-based reference regions, and (3) selecting optimal knot locations at maximum gradient points within each region. The resulting knot vector is then applied to SDD for accurate approximation of non-smooth and locally oscillating responses. A modal analysis of a lower control arm demonstrates that SDD with the proposed knot selection achieves higher accuracy than SDD with uniformly or randomly spaced knots, and also a Gaussian process surrogate model. The proposed SDD exhibits the lowest relative variance error (2.89%), compared to SDD with uniformly spaced knots (12.310%), randomly spaced knots (15.274%), and Gaussian process (5.319%) in the first natural frequency distribution. All surrogate models are constructed using the same 401 simulation datasets, and the relative errors are evaluated against a 2000-sample Monte Carlo simulation. The scalability and applicability of proposed method are demonstrated through stochastic and reliability analyses of mathematical functions ($N=1,3$) and a lower control arm system ($N=10$). The results confirm that both second-moment statistics and reliability estimates can be accurately achieved with only a few hundred function evaluations or finite element simulations.
@@
Uncertainty quantification, Surrogate model, Optimal knot vector, Spline dimensional decomposition, Dynamic system, Stochastic properties, Reliability analysis
@@ 1 Introduction
Uncertainty quantification (UQ) plays a pivotal role in the design and analysis of reliable engineering systems, as real-world input parameters—such as material properties, geometric dimensions, boundary conditions, and loading forces—are subject to inherent variability and randomness. Forward UQ propagates uncertainties from the input space to the output space. This process enables the estimation of key statistical properties such as the mean, variance, confidence intervals, and tail probabilities (e.g., failure probability) to ensure system safety and requirements under uncertain conditions. A common approach to forward UQ is the use of crude Monte Carlo simulation (MCS), wherein a large number of independent realizations of input variables are generated based on their prescribed probability distributions, and the corresponding system responses are computed. While MCS is conceptually straightforward and asymptotically accurate, it suffers from extremely slow convergence rates, particularly when estimating low-probability events or when the computational model is expensive to evaluate. Consequently, the method becomes computationally inefficient or even infeasible for high-fidelity engineering models [29, 9].
@@
Surrogate modeling has emerged as a powerful strategy to mitigate the high computational cost of UQ, especially in scenarios involving complex or high-fidelity simulation models. Surrogate models approximate the input-output relationship using a limited number of simulation samples, thereby enabling efficient propagation of uncertainty with significantly reduced computational effort [28, 17]. A range of state-of-the-art surrogate modeling techniques have been developed for UQ applications, including artificial neural networks (ANNs) [30], polynomial chaos expansion (PCE) [13], polynomial dimensional decomposition (PDD) [19], spline chaos expansion (SCE) [20], spline dimensional decomposition (SDD) [22], Kriging [10] and more [24, 15, 12]. Each method offers unique advantages in terms of scalability, accuracy, and suitability for various types of uncertainty and system behavior. Among these, SDD particularly excels at capturing strong nonlinearity and both local and global nonsmoothness in the response surface. SDD constructs the surrogate using hierarchical spline functions combined with ANOVA decomposition, allowing it to efficiently capture discontinuities and steep gradients in high-dimensional input spaces (e.g., greater than 20 dimensions).
@@
The SDD uses B-splines as basis functions in the surrogate model [20] to handle non-smooth regions. B-spline knot vectors divide the domain into intervals, applying polynomial functions of specific orders to capture localized variations precisely. This approach makes B-spline-based surrogate models especially effective for accurately predicting nonlinear responses—such as harmonic or transient dynamic responses—exhibiting nonsmooth or non-differentiable characteristics in dynamic engineering systems [21]. However, the accuracy of SDD relies heavily on the placement of knots as these affect the model’s ability to handle regions with sparse data or sharp gradients [7, 25]. As a result, the development of algorithms that adjust a proper placement of internal knots has enabled the construction of adaptive B-spline based models [23, 27, 6]. The work [23] proposes a modified Not-a-Knot condition that enforces zero second derivatives at boundary points for B-splines, incorporating hierarchical structures with spatial adaptivity to compute expansion coefficients efficiently. The work [27] proposes a Globally Convex Moving Mesh Algorithm-based knot vector optimization for Non-Uniform Rational B-Splines (NURBS) that uses an objective function combining data fitting and smoothing terms to reduce noise and prevent overfitting. The work [6] determines optimal knot vectors and expansion coefficients by minimizing an objective function based on prediction error and smoothness using evolutionary strategies and interior-point algorithms. A recent work [7] demonstrates high accuracy in numerical examples by employing sequential quadratic programming (SQP) to determine optimal knot vectors. These methods, however, often require iterative optimization procedures that repeat evaluations of objective functions, which can be computationally demanding, particularly when integrated into design optimization workflows or applied to high-dimensional UQ problems. A work [11] uses SDD for solving robust design optimization, but do not considers optimal knot vectors. Therefore, we need to develop a more computationally efficient and accurate method for determining optimal knot vectors.
@@
This work introduces a novel method for determining optimal knots in SDD for forward UQ in dynamic systems without optimization processes or additional training data. In the method, we interpolate a response profile using a sample dataset that is used for training SDD. We then define reference regions based on the specified number of internal knots. Within each reference region, we identify and set the point where the gradient of the interpolated profile reaches its maximum as an internal knot. We then use these internal knots to construct orthonormal spline basis functions for SDD via a whitening transformation. To estimate the expansion coefficients of SDD, we use both standard least squares (SLS) and least absolute shrinkage and selection operator (LASSO) regression, with the latter being particularly effective for high-dimensional problems where SLS may suffer from overfitting or ill-conditioning. This work is structured as follows. Section 2 provides the theoretical background, including the definitions of input and output random variables and an introduction to the stochastic dynamic problem. In addition, the summary of the SDD principle, second-moment properties, and failure probability. Section 3 proposes the determination of optimal knot vector and an SDD with the proposed knot vector. Also, convergence and stochastic properties are presented. Section 4 reports numerical examples. In Section 5, we offer conclusions and an outlook for future work.
@@ 2.4 Stochastic dynamic problem
Consider a linear, $M$-degree-of-freedom dynamic system with a stochastic mass matrix $\mathbf{M}(\mathbf{X})\in\mathbb{R}^{M\times M}$, a stochastic damping matrix $\mathbf{C}(\mathbf{X})\in\mathbb{R}^{M\times M}$, and a stochastic stiffness matrix $\mathbf{K}(\mathbf{X})\in\mathbb{R}^{M\times M}$. The mass matrix $\mathbf{M}(\mathbf{X})$ is real, symmetric, and positive-definite, while the stiffness and damping matrices, $\mathbf{K}(\mathbf{X})$ and $\mathbf{C}(\mathbf{X})$, are real, symmetric, and positive semi-definite. The damping matrix $\mathbf{C}(\mathbf{X})$ can be proportional or non-proportional; if proportional, it is expressed as a linear combination of the mass and stiffness matrices. When subjected to an $M$-dimensional deterministic force vector $\mathbf{f}(t)\in\mathbb{R}^{M}$, the system’s equation of motion in the time domain $t\in[0,T],\leavevmode\nobreak\ T\in\mathbb{R}^{+}$, is
@@ 2.6 Second-moment properties and failure probability via SDD
Consider the accuracy of SDD approximation, which heavily depends on the selection of the knot vector. Given a stochastic dynamic system and a computationally intensive model, the objective is to efficiently determine an optimal knot vector to enhance the accuracy of the SDD approximation for stochastic dynamic analysis.
@@ 3.2 Interpolation method for an optimal knot vector
In the proposed method, reference regions are introduced to address high-dimensional problems where the response shows complex oscillatory behavior. As illustrated in Figure 2, such behavior includes multiple local peaks. This makes it difficult to capture significant variation with a single representative knot. To overcome this limitation, multiple internal knots should be identified rather than relying on a single representative point. Instead of searching across the entire input domain at once, it is more effective to divide the domain into local regions and identify gradient maximum point within each region. Based on this idea, the input space is uniformly partitioned into reference regions $R_{k,j}$. This uniform division ensures that the entire domain is equally considered without additional complexity. The number of reference regions is set equal to the number of internal knots, allowing each internal knot to capture the most significant variation within its region by gradient method.
@@
Algorithm 1 outlines the implementation steps of the proposed interpolation-based knot selection procedure. We obtain input-output data samples $\{\mathbf{x}^{(l)},y(\mathbf{x}^{(l)})\}_{l=1}^{L}$ from either experimental measurements or numerical simulations. The response profile is estimated using linear interpolation based on the sampled input-output pairs. The input domain is then partitioned into $I_{k}$ reference regions $R_{k,j}$, i.e,
@@
$\tau_{k,j}=\operatorname*{arg\,max}_{x^{(l^{\prime})}\in R_{k,j}}\left|\frac{f(x^{(l^{\prime})})-f(x^{(l^{\prime}-1)})}{x^{(l^{\prime})}-x^{(l^{\prime}-1)}}\right|,\quad l^{\prime}=2,\ldots,L,\leavevmode\nobreak\ j=1,\dots,I_{k},$ (17)
Here, $L$ denotes the total number of input-output samples $\{\mathbf{x}^{(l)},y(\mathbf{x}^{(l)})\}_{l=1}^{L}$ used to estimate the SDD coefficients, as detailed in Section 3.3.3. This procedure does not require any additional computational cost, as it operates directly on existing samples. The procedure also avoids additional optimization steps, thereby enabling efficient identification of nonlinear behavior within each subdomain.
@@
3:Sort the input-output samples $\{\mathbf{x}^{(l)},y(\mathbf{x}^{(l)})\}_{l=1}^{L}$
4:Estimate output profile by a linear interpolation method, $l^{\prime}=2,\dots,L$, $k=1,\dots,N$,
$\quad f(x_{k})=f(x_{k}^{(l^{\prime}-1)})+\frac{f(x^{(l^{\prime})}_{k})-f(x_{k}^{(l^{\prime}-1)})}{x^{(l^{\prime})}_{k}-x_{k}^{(l^{\prime}-1)}}(x_{k}-x_{k}^{(l^{\prime}-1)}),\quad x_{k}^{l^{\prime}-1}\leq x_{k}\leq x_{k}^{l^{\prime}}.$
@@
$\tau_{k,j}=\operatorname*{arg\,max}_{x^{(l^{\prime})}\in R_{k,j}}\left|\frac{f(x^{(l^{\prime})})-f(x^{(l^{\prime}-1)})}{x^{(l^{\prime})}-x^{(l^{\prime}-1)}}\right|,\quad l^{\prime}=2,\ldots,L,\leavevmode\nobreak\ j=1,\dots,I_{k}.$
@@ 4 Numerical examples
Three numerical examples are presented to validate the proposed interpolation-based knot selection method for SDD approximation in solving UQ problems for stochastic dynamic analysis. The first example focuses on frequency response analysis, while the second and third address modal analysis. To ensure a fair comparison of accuracy across different knot vectors(uniformly spaced, randomly spaced, proposed method), the degree $p_{k}$, knot vector $\xi_{k}$, and the number of subintervals $I_{k}$ are kept consistent for all cases. Furthermore, all knot vectors are configured as $(p_{k}+1)-open$ knots. The proposed optimal knot vector method does not involve any additional optimization steps. As a result, the computational time required to construct the optimal knot vector is nearly identical to that of constructing uniform and random knot vectors. Therefore, this work does not include any time comparisons.
We additionally compare the proposed interpolation-based optimal knot selection method with a Gaussian Process Regression (GP) across all examples. For the output $y(\mathbf{x})$, we use GP as a benchmark surrogate model, defined as
$y(\mathbf{x})\sim\mathcal{GP}(m(\mathbf{x}),k(\mathbf{x},\mathbf{x}^{\prime})),$ (30)
where $m(\mathbf{x})$ is the mean function and $k(\mathbf{x},\mathbf{x}^{\prime})$ is the covariance function between two input vectors $\mathbf{x}$ and $\mathbf{x}^{\prime}$. Here, $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{N}$ are sampled realizations of the random vector $\mathbf{X}=(X_{1},\ldots,X_{N})^{\intercal}$. In this study, GP is implemented with a zero mean function, i.e., $m(\mathbf{x})=0$, and a Matern covariance kernel defined as
$k(\mathbf{x},\mathbf{x}^{\prime})=\sigma^{2}\cdot\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\|\mathbf{x}-\mathbf{x}^{\prime}\|}{\ell}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}\|\mathbf{x}-\mathbf{x}^{\prime}\|}{\ell}\right),$ (31)
where $\sigma^{2}\in\mathbb{R}_{0}^{+}$ is the signal variance, $\ell\in\mathbb{R}_{0}^{+}$ is the length scale, $\nu\in\mathbb{R}_{0}^{+}$ is the smoothness parameter, $\Gamma(\cdot)$ is the gamma function, and $K_{\nu}(\cdot)$ is the modified Bessel function of the second kind. The $\ell$, $\nu$) of the Matern kernel are empirically tuned to minimize the prediction error. The detailed values used in each case are provided in the corresponding example sections.

### v2

@@ Abstract
Forward uncertainty quantification in dynamical systems is challenging due to non-smooth or locally oscillating nonlinear behaviors. Spline dimensional decomposition (SDD) effectively addresses such nonlinearity by partitioning input coordinates via knot placement; yet, its accuracy is highly sensitive to the location of internal knots. We propose a computationally efficient, interpolation-based method for optimal knot selection in SDD. The method involves three steps: (1) interpolating input-output profiles, (2) defining subinterval-based reference regions, and (3) selecting optimal knot locations at maximum gradient points within each region. The resulting knot vector is then applied to SDD for accurate approximation of non-smooth and locally oscillating responses. A modal analysis of a lower control arm demonstrates that SDD with the proposed knot selection achieves higher accuracy than SDD with uniformly or randomly spaced knots, and also a Gaussian process surrogate model. In this example, the proposed SDD exhibits the lowest relative variance error (2.89%) in the first natural frequency distribution, compared to SDD with uniformly spaced knots (12.310%), randomly spaced knots (15.274%), and the Gaussian process (5.319%). The scalability and applicability of proposed method are demonstrated through stochastic and reliability analyses of one- and three-dimensional mathematical functions, along with a ten-dimensional lower control arm model. The results confirm that both second-moment statistics and reliability estimates can be accurately achieved with only a few hundred function evaluations or finite element simulations.
@@
Uncertainty quantification, Surrogate model, Optimal knot vector, Spline dimensional decomposition, Dynamical system, Stochastic properties, Reliability analysis
@@ 1 Introduction
Uncertainty quantification (UQ) plays a pivotal role in the design and analysis of reliable engineering systems, as real-world input parameters—such as material properties, geometric dimensions, boundary conditions, and loading forces—are subject to inherent variability and randomness. Forward UQ propagates uncertainties from the input space to the output space. This process enables the estimation of key statistical properties such as the mean, variance, confidence intervals, and tail probabilities (e.g., failure probability) to ensure system safety and requirements under uncertain conditions. A common approach to forward UQ is the use of crude Monte Carlo simulation (MCS), wherein a large number of independent realizations of input variables are generated based on their prescribed probability distributions, and the corresponding system responses are computed. While MCS is conceptually straightforward and asymptotically accurate, it suffers from extremely slow convergence rates, particularly when estimating low-probability events or when the computational model is expensive to evaluate. Consequently, the method becomes computationally inefficient or even infeasible when applied to high-fidelity engineering models [30, 10].
@@
Surrogate modeling has emerged as a powerful strategy to mitigate the high computational cost of UQ, especially in scenarios involving complex or high-fidelity simulation models. Surrogate models approximate the input-output relationship using a limited number of simulation samples, thereby enabling efficient propagation of uncertainty with significantly reduced computational effort [29, 18]. A range of state-of-the-art surrogate modeling techniques have been developed for UQ applications, including artificial neural networks (ANNs) [32], polynomial chaos expansion (PCE) [14], polynomial dimensional decomposition (PDD) [21], spline chaos expansion (SCE) [22], spline dimensional decomposition (SDD) [24], reduced order modeling [9], Kriging [11] and more [16, 13, 31]. Each method offers unique advantages in terms of scalability, accuracy, and suitability for various types of uncertainty and system behavior. Among these, SDD particularly excels at capturing strong nonlinearity and both local and global nonsmoothness in the response surface. SDD constructs the surrogate using hierarchical spline functions combined with ANOVA decomposition, allowing it to efficiently capture discontinuities and steep gradients in high-dimensional input spaces.
@@
The SDD uses B-splines as basis functions in the surrogate model [22] to handle non-smooth regions. B-spline knot vectors divide the domain into intervals, applying polynomial functions of specific orders to capture localized variations precisely. This approach makes B-spline-based surrogate models especially effective for accurately predicting nonlinear responses—such as harmonic or transient dynamic responses—exhibiting nonsmooth or non-differentiable characteristics in dynamic engineering systems [23]. However, the accuracy of SDD relies heavily on the placement of knots as these affect the model’s ability to handle regions with sparse data or sharp gradients [7, 26]. As a result, the development of algorithms that adjust a proper placement of internal knots has enabled the construction of adaptive B-spline based models [25, 28, 6]. A recent study [25] proposes a modified Not-a-Knot condition that enforces zero second derivatives at boundary points for B-splines, incorporating hierarchical structures with spatial adaptivity to compute expansion coefficients efficiently. Another recent study [28] proposes a Globally Convex Moving Mesh Algorithm-based knot vector optimization for Non-Uniform Rational B-Splines (NURBS) that uses an objective function combining data fitting and smoothing terms to reduce noise and prevent overfitting. An earlier study [6] determines optimal knot vectors and expansion coefficients by minimizing an objective function based on prediction error and smoothness using evolutionary strategies and interior-point algorithms. A separate study work [7] demonstrates high accuracy in numerical examples by employing sequential quadratic programming (SQP) to determine optimal knot vectors. These methods, however, often require iterative optimization procedures that repeat evaluations of objective functions, which can be computationally demanding, particularly when integrated into design optimization workflows or applied to high-dimensional UQ problems. Another study [12] uses SDD for solving robust design optimization, but does not consider optimal knot vectors. Therefore, we need to develop a more computationally efficient and accurate method for determining optimal knot vectors.
@@
This work introduces a novel method for determining optimal knots in SDD for forward UQ in dynamical systems without relying on optimization procedures or requiring additional training data. In the method, we interpolate a response profile using a sample dataset that is used for training SDD. We then define reference regions based on the specified number of internal knots. Within each reference region, we identify and set the point where the gradient of the interpolated profile reaches its maximum as an internal knot. We then use these internal knots to construct orthonormal spline basis functions for SDD via a whitening transformation. To estimate the expansion coefficients of SDD, we use both standard least squares (SLS) and least absolute shrinkage and selection operator (LASSO) regression, with the latter being particularly effective for high-dimensional problems where SLS may suffer from overfitting or ill-conditioning.
This paper is structured as follows. Section 2 provides the theoretical background, including the definitions of input and output random variables and an introduction to the stochastic dynamic problem. In addition, the summary of the SDD principle, second-moment properties, and failure probability. Section 3 proposes the determination of optimal knot vector and an SDD with the proposed knot vector. Also, convergence and stochastic properties are presented. Section 4 reports numerical examples. In Section 5, we offer conclusions and an outlook for future work.
@@ 2.4 Stochastic dynamic problem
Consider a linear, $M$-degree-of-freedom dynamical system with a stochastic mass matrix $\mathbf{M}(\mathbf{X})\in\mathbb{R}^{M\times M}$, a stochastic damping matrix $\mathbf{C}(\mathbf{X})\in\mathbb{R}^{M\times M}$, and a stochastic stiffness matrix $\mathbf{K}(\mathbf{X})\in\mathbb{R}^{M\times M}$. The mass matrix $\mathbf{M}(\mathbf{X})$ is real, symmetric, and positive-definite, while the stiffness and damping matrices, $\mathbf{K}(\mathbf{X})$ and $\mathbf{C}(\mathbf{X})$, are real, symmetric, and positive semi-definite. The damping matrix $\mathbf{C}(\mathbf{X})$ can be proportional or non-proportional; if proportional, it is expressed as a linear combination of the mass and stiffness matrices. When subjected to an $M$-dimensional deterministic force vector $\mathbf{f}(t)\in\mathbb{R}^{M}$, the system’s equation of motion in the time domain $t\in[0,T],\leavevmode\nobreak\ T\in\mathbb{R}^{+}$, is
@@ 2.6 Second-moment properties and failure probability via SDD
Consider the accuracy of SDD approximation, which heavily depends on the selection of the knot vector. Given a stochastic dynamical system and a computationally intensive model, the objective is to efficiently determine an optimal knot vector to enhance the accuracy of the SDD approximation for stochastic dynamic analysis.
@@ 3.2 Interpolation method for an optimal knot vector
In the proposed method, reference regions are introduced to address high-dimensional problems where the response shows complex oscillatory behavior. As illustrated in Figure 2, such behavior includes multiple local peaks. This makes it difficult to capture significant variation with a single representative knot. To overcome this limitation, multiple internal knots should be identified rather than relying on a single representative point. Instead of searching across the entire input domain at once, it is more effective to divide the domain into local regions and identify the maximum gradient point within each region. Based on this idea, the input space is uniformly partitioned into reference regions $R_{k,j}$. This uniform division ensures that the entire domain is equally considered without additional complexity. The number of reference regions is set equal to the number of internal knots, allowing each internal knot to capture the most significant variation within its region by gradient method.
@@
Algorithm 1 outlines the implementation steps of the proposed interpolation-based knot selection procedure. We obtain input-output data samples $\{\mathbf{x}^{(l)},y^{(l)}\}_{l=1}^{L}$, where $y^{(l)}=y(\mathbf{x}^{(l)})$, from either experimental measurements or numerical simulations. For the $k$th coordinate, we rearrange the dataset so that the input values are ordered as $x_{k}^{(l_{1})}<\cdots<x_{k}^{(l_{L})}$, with corresponding output values $y^{(l_{1})},\ldots,y^{(l_{L})}$, where $\{l_{1},\ldots,l_{L}\}$ is a permutation of the index set $\{1,\ldots,L\}$. The output profile is estimated using linear interpolation of the sampled input-output pairs, obtaining a linear function $f(x_{k})$, as detailed in Algorithm 1. The input domain is then partitioned into $I_{k}$ reference regions $R_{k,j}$, i.e,
@@
$\tau_{k,j}=\operatorname*{arg\,max}_{x_{k}^{(l_{i})}\in R_{k,j}}\left|\frac{f(x_{k}^{(l_{i})})-f(x_{k}^{(l_{i-1})})}{x_{k}^{(l_{i})}-x_{k}^{(l_{i-1})}}\right|,\quad i=2,\ldots,L,\leavevmode\nobreak\ j=1,\dots,I_{k},$ (17)
where $x_{k}^{(l_{i-1})}<x_{k}^{(l_{i})}$ and $l_{i}\in\{1,\ldots,L\}$. This knot selection procedure does not require any additional computational cost, as it operates directly on existing samples $\{\mathbf{x}^{(l)},y^{(l)}\}_{l=1}^{L}$ used for estimating the SDD coefficients (see Section 3.3.3). The procedure also avoids additional optimization steps, thereby enabling efficient identification of nonlinear behavior within each subdomain.
@@
3:For each $k$th coordinate, rearrange the input-output samples $\{\mathbf{x}^{(l)},y^{(l)}\}_{l=1}^{L}$ so that $x_{k}^{(l_{1})}<\cdots<x_{k}^{(l_{L})}$, with the corresponding outputs $y^{(l_{1})},\ldots,y^{(l_{L})}$.
4:Estimate output profile by a linear interpolation method, $i=2,\dots,L$, $k=1,\dots,N$,
$\quad f(x_{k})=y^{(l_{i-1})}+\frac{y^{(l_{i})}-y^{(l_{i-1})}}{x^{(l_{i})}_{k}-x_{k}^{(l_{i-1})}}(x_{k}-x_{k}^{(l_{i-1})}),\quad x_{k}^{(l_{i-1})}\leq x_{k}\leq x_{k}^{(l_{i})}.$
@@
$\tau_{k,j}=\operatorname*{arg\,max}_{x^{(l_{i})}\in R_{k,j}}\left|\frac{f(x^{(l_{i})})-f(x^{(l_{i-1})})}{x^{(l_{i})}-x^{(l_{i-1})}}\right|,\quad i=2,\ldots,L,\leavevmode\nobreak\ j=1,\dots,I_{k}.$
@@ 4 Numerical examples
Three numerical examples are presented to showcase the effectiveness of the proposed interpolation-based knot selection method for SDD approximation in solving UQ problems for stochastic dynamic analysis. The first example focuses on frequency response analysis, while the second and third address modal analysis. To ensure a fair comparison of accuracy across different knot vectors(uniformly spaced, randomly spaced, proposed method), the degree $p_{k}$, knot vector $\xi_{k}$, and the number of subintervals $I_{k}$ are kept consistent for all cases. Furthermore, all knot vectors are configured as $(p_{k}+1)-open$ knots. The proposed optimal knot vector method does not involve any additional optimization steps. As a result, the computational time required to construct the optimal knot vector is nearly identical to that of constructing uniform and random knot vectors. Therefore, this work does not include any time comparisons.
We additionally compare the proposed SDD with the Gaussian Process (GP) model across all examples. For the output $y(\mathbf{x})$, we use GP as a benchmark surrogate model, defined as $y(\mathbf{x})\sim\mathcal{GP}(m(\mathbf{x}),k(\mathbf{x},\mathbf{x}^{\prime})),$ where $m(\mathbf{x})$ is the mean function and $k(\mathbf{x},\mathbf{x}^{\prime})$ is the covariance function between two input vectors $\mathbf{x}$ and $\mathbf{x}^{\prime}$. Here, $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{N}$ are sampled realizations of the random vector $\mathbf{X}=(X_{1},\ldots,X_{N})^{\intercal}$. In this study, GP is implemented using the scikit-learn library [19], with a zero mean function, i.e., $m(\mathbf{x})=0$, and a Matérn kernel covariance kernel defined as $k(\mathbf{x},\mathbf{x}^{\prime})=\sigma^{2}\cdot\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\|\mathbf{x}-\mathbf{x}^{\prime}\|}{\ell}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}\|\mathbf{x}-\mathbf{x}^{\prime}\|}{\ell}\right),$ where $\sigma^{2}\in\mathbb{R}_{0}^{+}$ is the signal variance, $\ell\in\mathbb{R}_{0}^{+}$ is the length scale, $\nu\in\mathbb{R}_{0}^{+}$ is the smoothness parameter, $\Gamma(\cdot)$ is the gamma function, and $K_{\nu}(\cdot)$ is the modified Bessel function of the second kind. The hyperparameters ($\ell$, $\nu$) of the Matérn kernel are empirically tuned to minimize the prediction error. The detailed values used in each case are provided in the corresponding example sections. All numerical results were obtained using Python 3.12 on an Intel i5 13400 processor with 32 GB of RAM.